### v12

# Exact frequency-domain solutions for nonlinear pendulum-like dynamics

Abstract
Pendulum-like dynamics underlies a wide range of physical systems, from classical nonlinear oscillators to superconducting circuits and tunneling dynamics in cold-atom platforms. Here we present an exact frequency-domain formulation of the pendulum equation that applies uniformly across oscillatory, separatrix, and rotational regimes. The resulting spectral representation reveals a previously hidden unification: all regimes share the same analytic spectral structure and characteristic frequency scale, differing only by parity selection and the emergence of a continuous spectrum at the separatrix. This framework exposes regime transitions as symmetry-driven reorganizations in frequency space rather than changes in timescale, and identifies the stopping trajectory as a dynamical continuum limit reached without system-size scaling. Beyond providing closed-form solutions, the approach offers a transparent spectral insight into a broad class of classical and quantum pendulum-like systems.
Nonlinear dynamics; frequency-domain analysis; spectral structure; regime transitions; superconducting circuits; cold-atom systems
Introduction—Pendulum-like dynamics, as shown in Fig. 1, remains an active topic of research [1, 2, 3, 4, 5], with broad applications ranging from superconducting Josephson junctions [6, 7] and Bose–Einstein condensate tunneling [8, 9] to energy harvesting and artificial intelligence [10, 11, 12, 13]. For example, pendulum-based mechanisms can harvest energy from human motion [11] or ocean waves via piezoelectric coupling [12], while machine-learning models employ analytical pendulum solutions to recognize continuous nonlinear dynamics [13].
Figure 1: Pendulum-like dynamics across classical and quantum systems. (a) Superconducting Josephson junction, where the phase difference $\phi=\phi_{2}-\phi_{1}$ governs the supercurrent. (b) Neutral-atom (bosonic) Josephson junction with populations $N_{L}$ and $N_{R}$ and tunneling between condensates. (c) Nonlinear pendulum with angular displacement $\varphi$. All systems are described by the nonlinear phase equation $\ddot{\phi}+\Omega_{L}^{2}\sin\phi=0$, which admits swinging, stopping, and spinning regimes.
Energy-conserving pendulum-like dynamics are governed by the following equation:
$\frac{d^{2}}{dt^{2}}\phi(t)+\Omega_{L}^{2}\sin\phi=0.$ (1)
Here, $\phi(t)$ and $\Omega_{L}$ are the characteristic dynamical phase and the angular frequency of linear oscillation. $\Omega^{2}_{L}$ depends on physical parameters of the system. For example, $\Omega^{2}_{L}={g/L}$ in a classical pendulum [14, 15, 16], ${2eI_{c}}/({\hbar C})$ in an ideal Josephson junction [17], or $\frac{2J}{\hbar}\sqrt{\Lambda+\lambda}]^{2}$ in Bose–Einstein condensate tunneling [9].
However, despite its long history dating back to Galileo [18] and the foundational work of Euler and Jacobi, the nonlinear pendulum remains analytically incomplete in the frequency domain. While exact time-domain solutions exist in terms of the Jacobi elliptic sine function $\mathrm{sn}(z,k)$—whose exact Fourier series is known—their nested form obscures the spectral content of the motion [19, 20, 21]. For example, the exact solution takes the form $\phi(t)=2\arcsin[k\,\text{sn}(\Omega_{L}t+K,\,k)]$, so knowing the exact Fourier series of $\text{sn}(z,k)$ does not constitute an exact Fourier series of $\phi(t)$.
As a result, studies of the nonlinear pendulum have relied on numerical methods [22, 23] or sophisticated theoretical schemes [24, 25, 26] to compute the Fourier coefficients of $\phi(t)$, with increasingly complicated expressions for higher harmonics. Even in recent experimental studies of large-amplitude motion [27, 28], the theoretical coefficients are obtained by expanding $\sin\phi$ as a power series [29, 30]. Therefore, the Fourier coefficients are practically an approximation.
The absence of an exact closed-form expression for the Fourier coefficients underscores that frequency-domain analysis of the nonlinear pendulum has remained unresolved for centuries. Previous approaches have yielded only approximate or numerical results, limited to specific regimes or initial conditions. Here, we present complete and exact frequency-domain solutions for all three classes of motion—swinging, stopping, and spinning—establishing a unified analytic framework that resolves this long-standing problem.
Figure 2: Time- and frequency-domain representations of pendulum-like dynamics. (a) $\phi(t)$ for swinging ($\omega_{m}=\omega_{c}-10^{-4}$), stopping ($\omega_{m}=\omega_{c}$), and spinning ($\omega_{m}=\omega_{c}+10^{-4}$) motions, comparing the spectral solutions in Eqs. (3)–(6) with exact numerical results obtained from Jacobi elliptic functions. (b) Corresponding frequency-domain representation, showing that all three regimes share the same analytic spectral structure, while differing only by parity selection of harmonics (odd for swinging, even for spinning) and by a discrete-to-continuum transition at the separatrix.
Formalism—Nonlinear motions of an energy-conserving pendulum can be divided into three classes: Swinging, Stopping, and Spinning. These names are coined in this work to enhance intuitive interpretation. Consider the case where the pendulum starts at the bottom, $\phi_{0}=0$, with only $\omega_{0}$ to initiate the motion. Due to conservation of energy, the angular speed at the bottom is always maximal; thus, $\omega_{m}$ will be used to denote the angular speed at the bottom.
If $\omega_{m}$ that initiates the motion is too low, the pendulum will not reach the top and will inevitably fall back down. This is referred to as “Swinging” motion. If $\omega_{m}$ is just sufficient for the pendulum to reach the top, it will approach the top but never quite get there. This is because conservation of energy dictates that the closer the pendulum gets to the top, the slower it moves. Ideally and conceptually, the pendulum approaches the top asymptotically, and the angular speed decreases exponentially in time. This is referred to as “Stopping” motion. Finally, if $\omega_{m}$ is too large, the pendulum will continue to rotate in one direction. This is referred to as “Spinning” motion.
There exists a critical value of angular speed, $\omega_{c}$, which separates the motion into three distinct classes. Using conservation of energy, it is straightforward to show that
[…]
We now present the exact and unified frequency-domain formulation of the energy-conserving nonlinear pendulum:
swinging: $\phi(t)=\sum_{n\textbf{ odd}}c_{n}\sin(n\Omega_{0}t),$ (3)
stopping: $\phi(t)=2\arcsin\!\big[\tanh(\Omega_{L}t)\,\big],$ (4)
$\phi(t)=\int_{0}^{\infty}\!\!\!d\Omega\,C(\Omega)\sin(\Omega t),$ (5)
spinning: $\phi(t)=2\Omega_{0}t+\sum_{n\textbf{ even}}c_{n}\sin(n\Omega_{0}t).$ (6)
For simplicity, a closed-form solution for the stopping motion is shown in Eq. (4). For uniformity, however, an integral form can also be written as Eq. (5). In this form, all three classes of motion are now on the same footing, using the frequency domain.
The spectral coefficients $\{c_{n},C(\Omega)\}$ are:
swinging, spinning: $\quad\;c_{n}=\frac{4}{n\cosh(\kappa\frac{n\Omega_{0}}{\Omega_{L}})},$ (7)
stopping: $C(\Omega)=\frac{2}{\Omega\cosh(\kappa\frac{\Omega\;}{\Omega_{L}})},$ (8)
with the parameters $\kappa$, fundamental frequency $\Omega_{0}$, and period $T$ defined uniformly across all regimes as follows:
$\kappa\equiv K(\sqrt{1-k^{2}}),\quad\Omega_{0}=\frac{2\pi}{T},\quad T\equiv\frac{4\Re[K(k)]}{\Omega_{L}}.$ (9)
Here, K(k) is the complete elliptic integral of the first kind [31, 32]. Formally, $\kappa$ and $\Omega_{0}$ remain well-defined positive real quantities for the full range of modulus $k$, so that the definitions above apply to all classes of motion.
Validation—Figure 2(a) shows examples of $\phi(t)$ from this work. As clearly seen, the pendulum starts at the bottom and continues to rise up, but eventually branches off into three distinctive patterns. Quantitatively, the exact numerical values obtained using solutions in the form of Jacobi elliptic functions are also plotted in Fig. 2(a). The agreement is found to be within 14 significant digits, a typical limit of computer precision. Therefore, the solutions in this work are also exact. In addition, the spectral formalism presented here can be readily extended to arbitrary initial conditions by introducing a phase-shift parameter $\delta$, which shifts the time origin forward or backward (see Supplemental Material).
Traditionally, the perturbation method has been the primary theoretical tool for studying the frequency content of the nonlinear pendulum [29, 30]. However, its scope has been limited to a swinging motion where the pendulum is initially at rest, positioned at an amplitude $\phi_{0}$. As a result, the $\phi(t)$ is often written as a Fourier series of odd harmonics in the perturbation methods. For example, the recent experimental study [28] has used the following coefficients:
$A_{1}$ $=+\left(\phi_{0}+\frac{1}{192}\phi_{0}^{3}+\frac{17}{120\cdot 512}\phi_{0}^{5}+\cdots\right)$ (10)
$A_{3}$ $=-\left(\frac{1}{192}\phi_{0}^{3}+\frac{1}{6\cdot 512}\phi_{0}^{5}+\cdots\right)$
$A_{5}$ $=+\left(\frac{1}{40\cdot 512}\phi_{0}^{5}+\cdots\right)$
Figure 3: Convergence of frequency-domain solutions. Maximum error of the angular displacement $\phi(t)$ relative to exact numerical solutions, comparing the perturbation method with the exact spectral solution presented in this work. Results are shown for two initial amplitudes: $\phi_{0}=\pi/6$ (solid circles) and $\phi_{0}=\pi/2$ (open circles). The horizontal axis denotes the perturbation order $n$ for the perturbative approach, or equivalently the number of harmonics retained in the spectral solution. Both methods exhibit exponential convergence, but the closed-form coefficients in Eq. (7) achieve machine precision with fewer harmonics, particularly at larger amplitudes.
Figure 3 provides a comparative view of the convergence performance between the perturbation method and the solution presented in this work. Both methods exhibit rapid exponential convergence, as indicated by the straight line on the logarithmic scale. However, the swinging solution in Eq. (3) appears to converge faster as the number of harmonics increases.
Figure 4: Transition between regimes.
Discussion—Beyond classical physics, non-intuitive applications of the nonlinear pendulum are found in emerging fields of quantum physics, such as the superconducting Josephson junction [6, 7] and, most recently, Bose–Einstein condensate tunneling [8, 9], as schematically illustrated in Fig. 1.
For example, a Josephson junction (JJ) has been used extensively to construct qubits [33, 34, 35] in devices like IBM’s quantum processors, Google’s Sycamore Processor, and Zuchongzhi processors. JJ can be understood via BCS theory [36], where the key dynamical parameter is the phase difference $\phi$ between the wavefunctions of the two superconducting islands. This is directly analogous to the angle $\phi$ in the classical pendulum. Following standard circuit theory, the dynamical equation for an ideal JJ can be written as [37, 17]:
$\frac{d^{2}}{dt^{2}}\phi(t)+\frac{2eI_{c}}{\hbar C}\sin(\phi)=0$ (11)
Here, $C$ is the stray capacitance due to the junction’s geometrical arrangement, and $I_{c}$ is the critical tunneling current, which depends on the material, thickness, and area of the junction. As the phase difference $\phi$ fluctuates—analogous to the dynamics of the angle $\phi$ in the pendulum—so do the current through and the voltage across the junction. The current and the voltage are related to wavefunctions phase difference as follows: $I=I_{c}\sin\phi,V=\frac{\hbar}{2e}\frac{d\phi}{dt}$. In particular, the voltage across the junction plays similar role as the angular velocity of the nonlinear pendulum.
Another example of a quantum mechanical analogue involves two coupled Bose–Einstein condensates (BECs) of neutral bosonic atoms trapped in a double-well potential, known as a bosonic Josephson junction (BJJ) [38, 39, 40, 41]. The bosons can tunnel between the wells, and the key observable variables are the normalized atom number imbalance $n=\frac{N_{L}-N_{R}}{N_{L}+N_{R}}$ and the relative phase between the two condensates $\phi=\phi_{L}-\phi_{R}$. The dynamics of the BJJ, modeled as a pendulum, have been extensively studied in the notable work of Pigneur and Schmiedmayer [9] where the dynamical equation can be written as:
$\frac{d^{2}}{dt^{2}}\phi(t)+[\frac{2J}{\hbar}\sqrt{\Lambda+\lambda}]^{2}\sin(\phi)=0.$ (12)
Here, $\{J,\Lambda,\lambda\}$ are physical constants characterizing tunneling energy and inter-atomic interaction. Similar to JJ, where the voltage plays the role of angular velocity in classical pendulum, $n(t)$ in BJJ is related to the phase difference as $n(t)=\frac{\hbar}{2J}\frac{1}{\Lambda+\lambda}\frac{d\phi}{dt}$. Therefore, $n(t)$ is analogous to $\omega(t)$ in classical pendulum.
Figure 2(b) reveals that the distinction between swinging and spinning dynamics is encoded entirely in the parity structure of the frequency spectrum. Swinging motion occupies odd harmonics, while spinning occupies even harmonics with a linear phase drift. Crucially, aside from this parity selection rule, the two regimes share the same spectral structure: the harmonic amplitudes are governed by an identical functional form,
$c_{n}=\frac{4}{n\cosh\!\left(\kappa\frac{n\Omega_{0}}{\Omega_{L}}\right)}.$ (13)
In this sense, parity emerges as the organizing principle in frequency space: swinging and spinning are spectrally identical, distinguished solely by parity selection.
Figure 4(a) illustrates a central and nontrivial consequence of the exact spectral solutions: the transition between swinging and spinning dynamics occurs without any change in the fundamental spectral scale. As the system approaches the separatrix from either side, the period of motion diverges logarithmically with identical asymptotic behavior. In dimensionless time units,
$T_{\mathrm{swing}}\sim T_{\mathrm{spin}}\sim 4\ln\!\left(\frac{4}{\sqrt{\delta\omega}}\right),$ (14)
where $\delta\omega=|\omega_{m}-\omega_{c}|$ measures the distance from the critical point. As a result, the fundamental frequency $\Omega_{0}=2\pi/T$ vanishes in the same manner for both regimes, and the spectral decay scale $\kappa$ is preserved across the transition. This shows that the swinging–spinning transition is driven by a redistribution of spectral weight, while the underlying frequency structure remains unchanged.
Figure 4(b) highlights a discrete-to-continuum transition that is fundamentally different from those encountered in most areas of physics. In essentially all familiar examples, the emergence of a continuous spectrum is driven by an external thermodynamic or spatial limit. For phonons or photons, discrete modes become continuous only as the size of the confining cavity grows without bound. In electronic systems, discrete molecular energy levels evolve into continuous bands as the number of unit cells increases toward the bulk limit. Similarly, in quantum wells, bound states merge into a continuum as the spatial confinement is weakened or the system size expands. In each case, the continuum arises because the system itself is enlarged and acquires infinitely many degrees of freedom.
The pendulum-like dynamics studied here depart sharply from this paradigm. The system size is fixed, no additional degrees of freedom are introduced, and no spatial or thermodynamic limit is taken. Yet, as the separatrix is approached, the frequency spectrum becomes continuous while preserving the same spectral structure. This discrete-to-continuum transition is therefore not a consequence of system size, geometry, or dimensionality, but is instead generated purely by the nonlinear dynamics of a single degree of freedom.
This distinction is difficult to recognize in the time domain, where the stopping motion appears as a singular trajectory and the separatrix is often treated as a pathological boundary between qualitatively different behaviors. Only through the exact frequency-domain formulation developed here—where swinging, stopping, and spinning share a unified spectral structure—does it become possible to identify the stopping regime as the common continuum limit of the swinging and spinning spectra.
Conclusion—We have presented a complete and exact frequency-domain formulation of energy-conserving pendulum-like dynamics that unifies swinging, stopping, and spinning motions within a single analytic framework. By exposing the spectral structure common to all regimes, this approach reveals that qualitative changes in motion arise not from altered frequency scales or deformed spectra, but from systematic reorganizations of spectral weight—through parity selection and, at the separatrix, through a dynamical collapse of the fundamental frequency. These results resolve a long-standing gap in the frequency-domain description of the nonlinear pendulum and provide a transparent interpretation of regime transitions that has remained inaccessible in conventional time-domain treatments.
Acknowledgments—The author thanks the Potjanachaisak family for their hospitality during a summer research visit in Chiang Khruea, Sakon Nakhon Province, Thailand.
Appendix A End Matter
For $k>1$, the elliptic integral and its argument may become complex; nevertheless, both $\kappa$ and $\Re[K(k)]$ remain well-defined positive real quantities. For numerical implementation, it is often convenient to use standard transformation properties of $K$, yielding:
$k<1:$ $\kappa=K(\sqrt{1-k^{2}}),$ $T=\frac{4K(k)}{\Omega_{L}}$
$k=1:$ $\kappa=\frac{\pi}{2},$ $T\to\infty$
$k>1:$ $\kappa=\frac{1}{k}K(\sqrt{1-1/k^{2}}),$ $T=\frac{4K(1/k)}{k\Omega_{L}}$
A.1 Theoretical Method
The detailed derivations of the exact frequency-domain solutions are provided in Supplemental Material. Here we summarize the guiding principles: the unified frequency-domain structure becomes transparent only after three key perspective shifts.
  1. 1.
    Choice of initial condition. Instead of adopting the traditional formulation in which the pendulum is released from rest at an amplitude $\phi_{0}$, we consider the motion initiated at the bottom. This perspective naturally promotes the initial angular speed $\omega_{m}$ to a characteristic parameter, enabling continuous transitions between all three classes of motion.
  2. 2.
    Choice of dynamical variable. The challenge in finding the exact Fourier coefficients arises from the fact that the Jacobi elliptic function is embedded within the $\arcsin$ function. However, we have found that the $\arcsin[\cdots]$ dependence of $\phi(t)$ can be avoided if one initially focuses on the angular velocity $\omega(t)$. Once its spectral content is obtained, one gains access to the frequency-domain solution of $\phi(t)$ via simple integration.
  3. 3.
    Choice of period in the spinning regime. The motion can be decomposed exactly into a linear phase drift plus a periodic phase oscillation. The oscillatory component remains strictly periodic. Exploiting this periodicity, we adopt twice the primitive oscillation period as the fundamental period—an admissible spectral choice—which aligns the harmonic lattice of the spinning solution with that of the swinging solution. This step is not merely technical: it reveals that both regimes share an identical analytic coefficient structure and differ only by parity selection. It also permits the separatrix to be identified as the continuum limit common to both regimes.
Together, each perspective shift reveals the solution layer by layer—(i) disentangling regime classification, (ii) removing the $\arcsin$ obstruction, and (iii) exposing parity and the continuum limit—thereby unifying the exact spectral structure across all regimes of motion.
A.2 Identical Spectral Structure
At first glance, the continuum coefficient $C(\Omega)$ appears to differ from the discrete harmonic coefficient $c_{n}$ by a factor of two. However, $c_{n}$ represents a discrete harmonic amplitude, whereas $C(\Omega)$ is a spectral density in frequency space. A direct comparison is:
$c_{n}\;\;\longleftrightarrow\;\;C(\Omega)\,\Delta\Omega.$
Near the separatrix, where the harmonic spacing collapses, the discrete lattice satisfies $\Delta\Omega=2\Omega_{0}$ due to parity selection. The effective discrete amplitude is therefore $C(\Omega)\,\Delta\Omega$, which restores the same analytic coefficient structure as $c_{n}$.
The factor of two in $C(\Omega)$ is thus not a deviation from unification, but the precise normalization factor required to preserve it.